# Jun-Yan He1, Zhi-Qi Cheng2, Chenyang Li1, Jingdong Sun2 Qi He2, Wangmeng Xiang1, Hanyuan Chen1, Jin-Peng Lan1, Xianhui Lin1 Kang Zhu1

### v1

@@ Abstract
footnotetext: We also thank Chongyang Xiang (Southwest Jiaotong University), Huang Xiang (Southwest Jiaotong University), Wei Liu (University of Michigan, Ann Arbor), and Kaibo Wang (Peddie School) for their contributions. Qi He is also from Southwest Jiaotong University. Kaibo Wang’s contributions were part of his ISEF competition. The project was completed during their visit to CMU.
@@ 1 Introduction
  • MetaDesigner introduces a multi-agent system that integrates evaluation and optimization modules to facilitate the discovery and realization of personalized artistic text styles. The system’s meticulous hyperparameter selection process ensures the generation of artworks that align with individual preferences. The accessibility and popularity of the model are demonstrated by the over 400,000 visits it has received on.
@@
  • MetaDesigner showcases the authors’ commitment to supporting the artistic typography research community by developing a carefully curated dataset of artistic texts. The dataset contains over 5,000 multilingual images, spanning English, Chinese, Japanese, and Korean, covering a wide range of artistic styles and cultural elements. This extensive dataset serves as a valuable resource for researchers and practitioners to explore new techniques, evaluate existing methods, and push the boundaries of artistic text generation.
@@ 3.2 Glyph Designer Agent
Figure 3: The hierarchical model tree consists of multi-level sub-categories to facilitate finer granularity in ToT model selection.
@@ 4.1 Comparison with SOTA Methods
WordArt Synthesis Success. As illustrated in Figure 5, SD-XL fails to accurately render text, exhibiting incomplete support for English. The TextDiffuser series performs adequately in English but falls short in Chinese, Korean, and Japanese. Anytext delivers superior results but does not successfully handle Korean and Japanese. DALL-E3 is limited to English support only. In contrast, MetaDesigner demonstrates robust performance across all tested languages, successfully generating high-quality WordArt. More details are in the supplementary material.
@@ 5 Conclusion
In this work, we introduce MetaDesigner, a revolutionary framework for synthesizing artistic typography by leveraging the strengths of Large Language Models (LLMs) to drive a user-centric design paradigm. At the core of MetaDesigner is a multi-agent system comprising the Pipeline, Glyph, and Texture agents, which collectively enable customized WordArt creations, from semantic enhancements to the imposition of complex textures. By merging generative AI with artistic typography, MetaDesigner represents a significant advancement in the field, democratizing typography design and enriching the digital art domain. The system’s ability to understand user preferences and generate contextually relevant WordArt sets it apart from existing solutions, making it a valuable tool for both professionals and enthusiasts.
@@ Appendix B More Details of ToT Model selection
Figure 10: The hierarchical model tree consists of multi-level sub-categories to facilitate finer granularity in ToT model selection.
@@ B.1 ToT Model Selection Template
B.2 ToT Model Selection Case:
@@
Figure 11: Effect of Merging LoRAs with Different Weights. The upper case is generated by fusing the Chinese Bronzes and Sketch Style LoRAs; the rest is by fusing the Comic and Croissant Style LoRAs.
@@
Figure 14: A variety of exquisite examples are provided. Our method supports multiple languages (Chinese, English, Japanese, Korean, Arabic numerals, etc.) and accommodates different word counts. The generated images can be applied to scenarios such as e-commerce marketing, film and entertainment, and personalized social networking.
@@ Appendix C More Details of Image Evaluation
C.1 ToT-LoRA Evaluation Prompt:
@@
C.2 LLaVA Evaluation Prompt:
@@ Appendix D More Details of WordArt Dataset
Linguistic Diversity and Selection Criteria. In our user study, a carefully curated subset of 20 prompts was deployed to generate WordArt, leveraging the proposed comparison methods. This subset was chosen to represent a broad linguistic palette, including English, Chinese, Japanese, and Korean, to test the MetaDesigner’s capacity to handle multilingual input effectively. The selection criteria for these prompts were based on their ability to represent the thematic and linguistic diversity of the larger dataset, as well as their potential relevance to real-world applications in e-commerce, education, and digital content creation.
Statistics and Analysis. Statistically, the dataset is balanced across the five themes, with each theme comprising 30 prompts. This ensures an equitable representation of styles and thematic content, facilitating a comprehensive evaluation of the MetaDesigner’s performance. The prompts were selected based on criteria such as cultural significance, popularity in contemporary digital art, and relevance to potential commercial and educational applications.
Applicability and Use Cases. The versatility of our dataset is further demonstrated through an extensive selection of cases, highlighted in Figure 14. These examples include but are not limited to, cartoon text, Chinese surnames, and Chinese solar terms, showcasing the dataset’s ability to support a wide range of artistic and linguistic expressions. Such diversity not only validates the MetaDesigner’s efficacy in generating contextually and culturally relevant WordArt but also underscores the dataset’s broad applicability across various domains. From e-commerce promotions leveraging customized WordArt to enhance consumer engagement, to educational settings where multilingual WordArt facilitates language learning and cultural education, the dataset serves as a foundational resource for exploring the intersection of AI, art, and language.
@@
WordArt Dataset Analysis. We have conducted a series of numerical analyses on our WordArt dataset, as illustrated in Figure 15. We have generated over two million images in our "Image Plaza", encompassing various image ratios, word counts, and topics. Images in the dataset can be retrieved on our WordArt space at https://modelscope.cn/studios/WordArt/WordArt/summary. We have also conducted numerical analyses on user preferences, as depicted in Figure 16. To date, users have created more than one million images on our space. The majority of users show a preference for an image aspect ratio of 16:9, typically input text consisting of no more than five characters. Furthermore, users exhibit a marked preference for images that feature an alpha channel.

### v2

@@ Abstract
footnotetext: We also thank Chongyang Xiang and Xiang Huang (Southwest Jiaotong University), Wei Liu (University of Michigan, Ann Arbor), and Kaibo Wang (Peddie School) for their contributions. Kaibo Wang’s work was part of his ISEF competition project. This research was conducted in collaboration with Carnegie Mellon University.
@@ 1 Introduction
  • MetaDesigner introduces a multi-agent system that integrates evaluation and optimization modules to facilitate the discovery and realization of personalized artistic text styles. The system’s hyperparameter selection process ensures the generation of artworks that align with individual preferences. The accessibility and popularity of the model are demonstrated by the over 500,000 visits it has received on.
@@
  • MetaDesigner showcases our commitment to supporting the artistic typography research community by developing a carefully curated dataset of artistic texts. The dataset contains over 5,000 multilingual images, spanning English, Chinese, Japanese, and Korean, covering a wide range of artistic styles and cultural elements. This extensive dataset serves as a valuable resource for researchers and practitioners to explore new techniques, evaluate existing methods, and push the boundaries of artistic text generation.
@@ 3.2 Glyph Designer Agent
Figure 3: The hierarchical model tree structure facilitates fine-grained ToT (Tree of Thought) model selection. The complete tree structure is available in the supplementary materials.
@@ 4.1 Comparison with SOTA Methods
WordArt Synthesis Performance. As illustrated in Figure 5, SD-XL fails to accurately render text, exhibiting incomplete support for English. The TextDiffuser series performs adequately in English but falls short in Chinese, Korean, and Japanese. Anytext delivers superior results but does not successfully handle Korean and Japanese. DALL-E3 is limited to English support only. In contrast, MetaDesigner demonstrates robust performance across all tested languages, successfully generating high-quality WordArt. More details are in the supplementary material.
@@ 5 Conclusion
In this work, we introduce MetaDesigner, a revolutionary framework for synthesizing artistic typography by leveraging the strengths of Large Language Models (LLMs) to drive a user-centric design paradigm. At the core of MetaDesigner is a multi-agent system comprising the Pipeline, Glyph, and Texture agents, which collectively enable customized WordArt creations, from semantic enhancements to the imposition of complex textures. By merging generative AI with artistic typography, MetaDesigner represents a significant advancement in the field, democratizing typography design and enriching the digital art domain. The system’s ability to understand user preferences and generate contextually relevant WordArt sets it apart from existing solutions, making it a valuable tool for both professionals and enthusiasts. Looking ahead, we envision MetaDesigner as a catalyst for further innovation in the intersection of AI and artistic expression. Future work could explore expanding the system’s capabilities to include more complex typographic manipulations, integration with 3D design tools, or adaptation to dynamic, real-time typography applications.
@@ Appendix B More Details of ToT Model selection
Figure 10: Hierarchical model tree with multi-level sub-categories enabling fine-grained ToT (Tree of Thought) model selection for diverse WordArt generation.
@@ B.1 ToT Model Selection Template
B.2 ToT Model Selection Case
@@
Figure 11: Impact of LoRA fusion with varying weights. Top: Merging Chinese Bronzes and Sketch Style LoRAs. Bottom: Combining Comic and Croissant Style LoRAs.
@@
Figure 14: Diverse examples of WordArt generated by our method. Supports multiple languages (Chinese, English, Japanese, Korean, Arabic numerals, etc.) and various word counts. Applications span e-commerce marketing, film and entertainment, personalized social networking, and more. The method adapts to different styles, textures, and themes, showcasing versatility across cultural and linguistic contexts. Enlarge for detailed viewing of intricate designs and multilingual rendering.
@@ Appendix C More Details of Image Evaluation
C.1 ToT-LoRA Evaluation Prompt
@@
C.2 LLaVA Evaluation Prompt
@@ Appendix D More Details of WordArt Dataset
Multilingual Scope and Selection Criteria. For our user study, we employed a curated subset of 20 prompts to generate WordArt using the proposed comparison methods. This subset was designed to showcase a rich linguistic tapestry, encompassing English, Chinese, Japanese, and Korean, thereby rigorously evaluating MetaDesigner’s adeptness in processing and rendering multilingual inputs. The selection process adhered to a multifaceted set of criteria, including each prompt’s capacity to represent the thematic and linguistic diversity of the broader dataset, its potential relevance to real-world applications across domains such as e-commerce, education, digital marketing, and creative content production, and its ability to challenge and showcase the system’s multilingual capabilities. Additionally, we considered each prompt’s cultural resonance and significance within diverse global contexts, ensuring an inclusive approach. This carefully crafted selection enables a comprehensive and nuanced assessment of MetaDesigner’s performance across various linguistic and cultural landscapes, providing valuable insights into its versatility and adaptability.
Dataset Composition and Thematic Balance. The dataset exhibits a balanced distribution across the 5 themes, with each theme, with each theme precisely encompassing 30 carefully crafted prompts. This equitable representation ensures a diverse range of styles and thematic content, enabling a thorough assessment of MetaDesigner’s capabilities. Prompt selection was guided by multiple criteria, including cultural significance, prominence in contemporary digital art, and relevance to potential commercial and educational applications. This careful curation process enhances the dataset’s utility for evaluating AI-driven artistic typography across various contexts and use cases.
Applicability and Use Cases. The dataset’s versatility is exemplified through a comprehensive array of cases, as illustrated in Figure 14. These examples span a diverse range, including cartoon text, Chinese surnames, and Chinese solar terms, demonstrating the dataset’s capacity to support a wide spectrum of artistic and linguistic expressions. This diversity not only validates MetaDesigner’s proficiency in generating contextually and culturally relevant WordArt but also underscores the dataset’s extensive applicability across various domains. The potential applications are far-reaching: from enhancing consumer engagement in e-commerce through customized WordArt promotions, to facilitating language learning and cultural education via multilingual WordArt in educational settings.
@@
Quantitative Analysis and User Preferences. Our comprehensive analysis of the WordArt dataset reveals significant insights, as illustrated in Figure 15. The "Image Plaza" now houses over two million images, showcasing a diverse range of image ratios, word counts, and thematic content. Interested parties can access and explore these images through our WordArt space at https://modelscope.cn/studios/WordArt/WordArt/summary. Additionally, we’ve conducted an in-depth examination of user preferences, as depicted in Figure 16. To date, our platform has facilitated the creation of more than one million user-generated images. The data indicates a strong user preference for the 16:9 image aspect ratio, with most inputs consisting of five characters or fewer. Notably, users demonstrate a marked inclination towards images featuring an alpha channel, highlighting a preference for designs with transparency or layering capabilities.